OpenStudy (anonymous):

What is the factored form of the expression? 12d2 + 4d – 1

OpenStudy (anonymous):

12(d2+1/3d+1/12) = 12(d2 + 2*1/6d + (1/6)^2) -12*(1/6)^2 -1 =12(d-1/6)^2 - 4/3 = 12[(d-1/6)^2 - 4/36] = 12[(d-1/6)^2 -(2/6)^2] = 12(d-1/6-2/6)(d-1/6+2/6) = 12(d-1/2)(d+1/6)